# Text-to-Image Rectified Flow as Plug-and-Play Priors

### v2

@@ 3.3 RFDS-Rev
To enhance the baseline RFDS, we draw inspiration from the learning mechanisms of rectified flows. It is important to recall that rectified flow essentially learns an optimal transport from the original noise distribution to the image distribution using straight flows [20]. The mode-seeking nature of the original RFDS, which scans different noises, leads to an averaged velocity and consequently blurred images. As shown on the left side of Fig. 2, when two noise points are sampled, although both give gradients towards the high-density areas in the image domain, the averaged gradient does not necessarily point in the correct direction. To address this issue, we first introduce a critical assumption:
Assumption 1: Given a sample $X_{1}$ from data distribution $\pi_{1}$, there is at most one corresponding $X_{0}$ from noise distribution $\pi_{0}$ that is coupled with $X_{1}$.
This assumption is straightforward, as the training process for rectified flow enforces a linear, non-crossing flow [20]. Based on this assumption, we propose RFDS-Reversal (RFDS-Rev), a two-stage method to improve the baseline RFDS. Specifically, as depicted on the right in Fig. 2, the first stage performs noise reversal. In the first stage, we use iRFDS to optimize each randomly sampled noise such that their reversal occupies a relatively static position within the noise distribution. Once the reversal is determined, in the second stage, we apply RFDS to the reversed noise. The algorithm of RFDS-Rev is summarized in Algorithm 1. In experiments, we observe that running iRFDS for just one step yields very satisfactory results. Unless specified otherwise, the experiments involving RFDS-Rev employ $n=1$ as the default configuration.
1 Initialize the learnable parameter $\theta$
@@ 4       Sample random timestep $t$
5       Sample random noise $\epsilon$
@@
7             Freeze $\theta$, optimize $\epsilon$ with iRFDS (Eq. 8)
@@
10      Optimize $\theta$ with optimized $\epsilon$ based on RFDS (Eq. 6)
@@
3.4 Classifier Free Guidance
The above equations are derived using the rectified flow network $v_{\phi}$. In practice, since the base rectified flow model [23] is trained with classifier free guidance [25], we observe that using a classifier free guidance version of the network $\hat{v}_{\phi}$ help to improve the performance. Ablation experiments of this problem are included in the Appendix.
@@ 4 Experiments
In this section, we conduct extensive experiments to compare our proposed methods with diffusion-based methods. Firstly, we focus on evaluating RFDS and RFDS-Rev in both 2D and 3D text-guided generation scenarios. Subsequently, we explore the performance of iRFDS in image inversion tasks. Unless otherwise noted, our rectified flow experiments utilize the large-scale text-to-image model InstaFlow [23]. For comparison, the diffusion-based methods are implemented using Stable Diffusion v2.1 [4] and Threestudio codebase [26]. In terms of base model generation FID, the rectified flow based model InstaFlow [23] that we used is slightly weaker compared to Stable Diffusion. Finally, we present qualitative results by applying our methods to the recently released Stable Diffusion 3 model [24]. We include the implementation details in the Appendix for further reference.
Figure 3: 2D Comparison. Text Prompts: A castle next to a mountain. An astronaut is riding a horse. A monster truck.
Figure 4: 3D Comparison. Text Prompts: A classic leatherette radio with dials, An antique wooden rocking horse, A fragrant pine Christmas wreath, A paint-splattered easel, A vibrant sunflower with green leaves, Hot popcorn jump out from the red striped popcorn maker.
4.1 RFDS vs. RFDS-Rev vs. Diffusion Priors
Toy Experiments on Optimization of 2D Case. We first conduct an intuitive experiment in the 2D setting by setting $x=\theta$. This setup removes the complexity of the 3D model, thereby showcasing the potential upper limit of 3D generation quality. We compare two of our proposed methods, the RFDS baseline and RFDS-Rev, against diffusion-based methods. Results are depicted in Fig. 3. We observe that our proposed baseline method, RFDS, achieves performance comparable to that of the diffusion-based SDS loss. With just one additional forward pass, our RFDS-Rev method significantly enhances the generation quality over the baseline RFDS. RFDS-Rev manages to produce very competitive results compared to VSD.
@@
Method - Dreamfusion (SDS) [14] ProlificDreamer (VSD) [15] RFDS RFDS-Rev
Base Model - Diffusion Diffusion Rectified Flow Rectified Flow
[…]
@@
Text-to-3D Generation by Lifting 2D Models. An important application of rectified flow or diffusion priors is lifting models from 2D to 3D. We present both quantitative and qualitative results for text-to-3D generation. The qualitative results are displayed in Fig. 4. We observe that both of our proposed methods are capable of generating high-quality 3D objects, with the RFDS-Rev method enhancing the baseline RFDS in terms of object detail and color accuracy. Qualitatively, the RFDS baseline shows competitive performance than the SDS loss used in diffusion-based methods. Compared to VSD, RFDS-Rev achieves a comparable level of detail in generated objects. Notably, VSD experiences issues with “exploding” artifacts, as highlighted in the red box in our figures. In contrast, the RFDS-Rev loss does not exhibit this problem. Additional qualitative results of RFDS-Rev are included in the Appendix.
@@
We also conduct quantitative experiments on text-to-3D benchmark T3Bench [27]. Results are demonstrated in Table 1. We observe that our proposed methods significantly outperform diffusion-based methods. Furthermore, the RFDS-Rev method sets a new state-of-the-art performance among all 2D lifting methods in this benchmark.
4.2 iRFDS vs. Rectified Flow Time Inversion vs. Diffusion Methods
We also study the performance of iRFDS in 2D image inversion and subsequent text-guided editing. Notably, a standard text-to-image rectified flow inherently supports the image-to-noise conversion due to its time-symmetric property. This can be achieved by simply reversing the flow from $t\rightarrow 1-0$ and taking a negative velocity, a method we refer to as “Rectified Flow Time Inversion”. Unlike “Rectified Flow Time Inversion”, which is essentially the inverse of the generation process, our proposed iRFDS treats the noise as an optimization objective and uses the pretrained model as a loss function to optimize the noise using optimizers like SGD.
@@
Image inversion is also a widely explored topic in diffusion-based methods, known for its capability to edit real-world images. We compare iRFDS against the popular method – the null inversion method on real-world images. Additionally, we evaluate our method against the diffusion prior based optimization method, the DDS loss [17], which directly optimizes the original images.
@@
The results of the inversion and editing are summarized in Fig. 5. We observe that “Rectified Flow Time Inversion” struggles to perform accurate inversions on real images. Compared to the diffusion-based null inversion, our proposed iRFDS method achieves very competitive performance. Meanwhile, the diffusion prior DDS loss tends to produce images with over-saturated colors. More visualization results can be found in Appendix.
@@
To further validate the performance improvements, we conducted quantitative experiments to calculate the CLIP score between the target caption and the image after editing. Additionally, we carry out a user study to evaluate the editing performance from the users’ perspective. The experiments are performed on 15 real images collected from the internet, with each image being edited three times using different captions. The results are summarized in Table. 2. We observe that our proposed method iRFDS achieves the best performance in terms of CLIP scores and user preference.
@@
Method DDS [17] Null Inversion [28] RF Time Inversion [20] iRFDS (Ours)
CLIP Score 0.294 0.298 0.268 0.306
User Preference 12.8% 31.7% 0.0% 55.5%
@@ 4.3 Ablation Experiments
iRFDS optimization steps in RFDS-Rev (Algorithm 1). As demonstrated in Fig. 6, we show results from both the 2D case and 3D case. We observe that performing the iRFDS for 1 step achieves the best efficiency and performance trade-off.
@@
w/ Network Jacobian vs. w/o Network Jacobian. As illustrated in Fig. 7, we observe that keeping the network Jacobian can hardly generate meaningful images even at 2D case.
Figure 7: Ablation experiments of ignoring the Rf network Jacobian. We observe that it is difficult to generate images by keeping the Jacobian of the Rf network.
Additional Ablation Results. We also include additional ablation results in the Appendix. The experiments include: ablation on the effect of CFG, RFDS-Rev vs. directly applying the same idea of VSD to rectified flow and a comparison of computational cost between our methods and diffusion-based methods.
4.4 Results on Stable Diffusion 3
Stable Diffusion 3 [24] is a new state-of-the-art text-to-image rectified flow model, trained with visual transformers and more robust text encoders. We observe that our proposed methods can also be applied to this model, resulting in significant performance improvements. The results of text-to-3D generation using Stable Diffusion 3 are presented in Fig. 8.
Figure 8: Text-to-3D generation with Stable Diffusion 3 and RFDS-Rev. By utilizing Stable Diffusion 3, our method can generate 3D assets with stunning object details.
5 Discussion
Advantages over diffusion based methods. Besides the performance improvements, we observe that methods based on rectified flow achieve faster convergence during the optimization of 3D models. Intuitively, if we rearrange Eq. 6, $x$ can be regarded as learning towards a direct objective $v_{\phi}+\epsilon$. In contrast, diffusion models rely on the difference in predicted noises to guide the optimization process, which is more circuitous. A detailed visual comparison is included in the Appendix.
@@ 6 Conclusion, Limitations and Broader Impacts
Limitations. Since the 2D model lacks training with camera pose information, our proposed methods encounter similar issues to the SDS and VSD losses in 3D generation, such as the multi-face issue. These limitations can be addressed in future studies by training pose-aware models using multi-view data, similar to current diffusion based approaches [29; 30; 31; 32; 33].
@@
Broader Impacts. Our approach significantly expands the potential applications of pretrained rectified flow models. We demonstrate that a text-to-image model can be adapted for both image editing and text-to-3D generation tasks. However, since we utilize pretrained text-to-image rectified flow models as our foundational network, our methods might also inherit the biases present in these base networks. For example, a model may predominantly depict professionals such as doctors or CEOs as males, thereby perpetuating existing societal stereotypes.
Appendix A Appendix
A.1 Implementation Details
@@
RFDS and RFDS-Rev for 3D generation. We use the large-scale rectified flow based text-to-image model InstaFlow [23] and the 3D model implicit model Instant-NGP [34]. Each 3D model is optimized for 15000 steps. We use a CFG of 50 for all 3D experiments and 2D toy experiments. The model is optimized with a resolution of 256 for the first 5000 steps and then 500 for the final 10000 steps. The experiments are carried out on NVIDIA A6000 GPUs. Generating one 3D scene using RFDS-Rev takes around 30 mins wall-clock time. We use $w(t)=1$ in all experiments.
@@
iRFDS for image inversion and editing. The inversion starts from a randomly sampled Gaussian noise. We optimize the noise for 1000 steps using iRFDS and CFG 1 with a learning rate of 3*10-3. To facilitate effective noise optimization, we add one additional loss to enforce the noise follows a Gaussian distribution. Specifically, we add a loss to enforce the mean and variance of the current noise to be zero and one respectively. After the noise is optimized, we change the caption to the target caption and run the forward flow for 5 steps using CFG 1.5. We use $w(t)=1$ in all experiments.
A.2 Additional Related Works

### v3

@@ 3.3 RFDS-Rev
1 Initialize the learnable parameter $\bm{\theta}$
@@ 4       Sample random timestep $t$
5       Sample random noise $\bm{\epsilon}$
@@
7             Freeze $\bm{\theta}$, optimize $\bm{\epsilon}$ with iRFDS (Eq. 10)
@@
10      Optimize $\bm{\theta}$ with optimized $\bm{\epsilon}$ based on RFDS (Eq. 8)
@@
To enhance the baseline RFDS, we draw inspiration from the learning mechanisms of Reflow (Liu et al., 2022; 2024). Recall that Reflow straightens the trajectory from noise to image, leading to one-to-one matching of the image and noise. In contrast, the mode-seeking nature of the original RFDS, which optimizes the image based on random noises, leads to an averaged velocity and consequently blurred images. As shown on the left side of Fig. 2, when two noise points are sampled, although both give gradients towards the high-density areas in the image domain, the averaged gradient does not necessarily point in the correct direction. To address this issue, we first introduce a critical assumption:
Assumption 1: For a model finetuned with Reflow, given a sample $\bm{x_{*}}$ from data distribution $p(\bm{x})$, there is at most one corresponding $\bm{\epsilon}$ from noise distribution $\mathcal{N}(0,\bm{I})$ that corresponds to $\bm{x_{*}}$.
This assumption is straightforward, as the training process for Reflow enforces a linear, non-crossing flow (Liu et al., 2022). Based on this assumption, we propose RFDS-Reversal (RFDS-Rev), a two-stage method to improve the baseline RFDS. Specifically, as depicted on the right in Fig. 2, the first stage performs noise reversal. In this stage, we use iRFDS to perform image inversion and optimize each randomly sampled noise such that their reversal occupies a relatively static position within the noise distribution. Once the reversal is determined, in the second stage, we apply RFDS to the reversed noise. The algorithm of RFDS-Rev is summarized in Algorithm 1. In experiments, we observe that running iRFDS for just one step yields very satisfactory results. Unless specified otherwise, the experiments involving RFDS-Rev employ $n=1$ as the default configuration. Although RFDS-Rev was initially designed with the assumption of models fine-tuned using Reflow, our experimental results reveal that it enhances the performance of the baseline RFDS on models both with and without Reflow training.
@@
3.4 Applying iRFDS and RFDS-Rev to Diffusion Models
As discussed in previous works (Song et al., 2020; Albergo et al., 2023), diffusion models trained with score-matching objectives have a deterministic PF-ODE form and can be expressed in the form of Eq. 2. We observe that by converting the score function to a velocity field, all three of our proposed methods become applicable to diffusion models. Additionally, we prove that our proposed RFDS baseline is identical to the vanilla SDS loss when applied to diffusion models. We include a detailed proof in Appendix Sec A.
3.5 Classifier Free Guidance (CFG)
The above equations are derived using the rectified flow network $\bm{v_{\phi}}$. In practice, since the base rectified flow models (Liu et al., 2024; Esser et al., 2024) are usually trained with classifier free guidance (Ho & Salimans, 2022), we observe that using a classifier free guidance version of the network $\bm{\hat{v}_{\phi}}$ helps to improve the performance. Ablation experiments on this problem are included in Appendix Sec E.
@@ 4 Experiments
In this section, we conduct extensive experiments to compare our proposed methods with diffusion-based approaches. We perform experiments on two types of rectified flow-based methods: Stable Diffusion v3 (SD3), trained with flow-matching but without Reflow finetuning, and InstaFlow, finetuned with Reflow. In terms of 2D performance, Stable Diffusion v3 is one of the state-of-the-art text-to-image models, trained with more data and using larger backbones. InstaFlow achieves results similar to those of Stable Diffusion v2.1 (SD2.1) (Rombach et al., 2022). First, we focus on evaluating RFDS and RFDS-Rev in both 2D and 3D text-guided generation scenarios. Subsequently, we explore the performance of iRFDS in image inversion tasks. The diffusion-based methods are implemented using SD2.1 and the Threestudio codebase (Guo et al., 2023a). We include the implementation details in the Appendix for further reference.
Figure 3: 2D Comparison. Text Prompts: “A castle next to a mountain”, “An astronaut is riding a horse” and “A monster truck”.
Figure 4: 3D Comparison. Text Prompts: “An antique wooden rocking horse”, “A bear holding a sign that says Hello world” and “Hot popcorn jump out from the red striped popcorn maker”.
4.1 RFDS vs. RFDS-Rev vs. Diffusion Priors vs. Diffusion RFDS-Rev
Toy Experiments on Optimization of 2D Case. We begin by conducting an intuitive experiment in a 2D setting, where we set $\bm{x}=\bm{\theta}$. This simplified setup removes the complexities of the 3D model, allowing us to demonstrate the potential upper limit of 3D generation quality. Results are shown in Fig. 3. Results reveal that our RFDS baseline, when integrated with InstaFlow, performs comparably to the diffusion-based SDS loss. Moreover, when combined with Stable Diffusion v3, the RFDS baseline significantly enhances generation quality.
We also apply our proposed method, RFDS-Rev, to two rectified flow-based models and the diffusion model Stable Diffusion v2.1, as discussed in Sec 3.4. To better approximate the real use cases, we use only one step of iRFDS to improve efficiency for all experiments. Our results indicate that RFDS-Rev significantly enhances generation performance on InstaFlow, moderately improves performance on SD3, and slightly improves performance on the diffusion model. These findings are consistent with the assumptions made when proposing the RFDS-Rev method. The Reflow process straightens the noise-to-data trajectory, making our assumptions particularly applicable to models finetuned with Reflow. Although SD3 is not finetuned with Reflow, our method still improves generation quality. However, in score-matching models like diffusion, noise is applied stochastically to corrupt the data, making it challenging to trace the original noise with just one step of iRFDS. For this reason, in the 3D experiments, we apply our methods only to flow-matching based approaches, and compare them with SDS-like diffusion-based approaches.
@@
Method Dreamfusion (SDS) ProlificDreamer (VSD) RFDS RFDS-Rev RFDS RFDS-Rev
Base Model Diffusion Diffusion InstaFlow InstaFlow SD3 SD3
Single Object 24.9 51.1 46.9 57.6 49.4 59.8
Surroundings 19.3 42.5 34.5 44.6 39.3 54.5
Multiple Objects 17.3 45.7 28.3 44.6 42.9 55.2
Average 20.5 46.4 36.5 48.9 43.9 56.5
@@
Text-to-3D Generation by Lifting 2D Models. An important application of rectified flow or diffusion priors is lifting models from 2D to 3D. We present both quantitative and qualitative results for text-to-3D generation. The qualitative results are displayed in Fig. 4. We observe that both of our proposed methods are capable of generating high-quality 3D objects, with the RFDS-Rev method enhancing the baseline RFDS in terms of object detail and color accuracy. Qualitatively, when applied to similar quality base models (InstaFlow and SD2.1), the RFDS baseline shows competitive performance compared to the SDS loss used in diffusion-based methods. Compared to VSD, RFDS-Rev achieves a comparable level of detail in generated objects. Notably, VSD experiences issues with “exploding” artifacts. In contrast, the RFDS-Rev loss does not exhibit this problem. When we apply our method to the more powerful SD3 base model, we observe that it outperforms diffusion priors. Moreover, RFDS-Rev further enhances the generation quality compared to the RFDS baseline on SD3. Additional qualitative results of RFDS-Rev are included in Appendix I.
@@
We also conduct quantitative experiments on the text-to-3D benchmark T3Bench (He et al., 2023). The dataset contains 300 text prompts for text-to-3D generation, making it the largest text-to-3D benchmark available. Results are demonstrated in Table 1. We report the quality score, derived by rendering multi-view images and calculating the CLIP (Radford et al., 2021) and ImageReward (Xu et al., 2024) scores between the rendered images and the corresponding text prompt. We observe that our proposed methods significantly outperform diffusion-based methods. Furthermore, the RFDS-Rev method sets a new state-of-the-art performance among all 2D lifting methods in this benchmark.
4.2 iRFDS vs. Diffusion Methods
@@
In this section, we study the performance of iRFDS in 2D image inversion and subsequent text-guided editing. Image inversion is also a widely explored topic in diffusion-based methods, known for its capability to edit real-world images. We compare iRFDS against a popular method – null inversion – on real-world images. Additionally, we evaluate our method against the diffusion prior based optimization method, the DDS loss (Hertz et al., 2023), which directly optimizes the original images.
@@
The results of the inversion and editing experiments are summarized in Fig. 5. Compared to the diffusion-based null inversion, our proposed iRFDS method delivers highly competitive performance. Additionally, the diffusion prior DDS loss tends to generate images with overly saturated colors. Among the different variants of our iRFDS methods, we find that the model finetuned with Reflow produces the best results. Further visualizations can be found in Appendix Sec. H.
@@
To further validate the performance improvements, we conducted quantitative experiments to calculate the CLIP score between the target caption and the image after editing. Additionally, we carry out a user study to evaluate the editing performance from the users’ perspective. The experiments are performed on 15 real images randomly collected from the internet, with each image being edited three times using different captions. The results are summarized in Table. 2. We observe that our proposed method iRFDS + InstaFlow achieves the best performance in terms of CLIP scores and user preference.
@@
Method DDS Null Inversion iRFDS (InstaFlow) iRFDS (SD3) iRFDS (Diffusion)
CLIP Score 0.294 0.298 0.306 0.297 0.296
User Preference 6.8% 18.6% 45.3% 17.1% 12.2%
@@ 4.3 Ablation Experiments
iRFDS optimization steps in RFDS-Rev (Algorithm 1). Fig. 6 shows results from both the 2D and 3D cases. We observe that performing the iRFDS for 1 step achieves the best efficiency and performance trade-off.
@@
w/ Network Jacobian vs. w/o Network Jacobian. As illustrated in Fig. 7, we observe that keeping the network Jacobian can hardly generate meaningful images even in the 2D case. This observation is consistent for both SD3 and InstaFlow.
Figure 7: Ablation experiments of ignoring the rectified flow network Jacobian. We observe that it is difficult to generate images by keeping the Jacobian of the rectified flow network.
Additional Ablation Results. We also include additional ablation results in Appendix Sec E. The experiments include: ablation on the effect of CFG, RFDS-Rev vs. directly applying the same idea of VSD to rectified flow and a comparison of computational cost between our methods and diffusion-based methods.
5 Summary of observations and Discussion
Summary of observations. 1), RFDS and RFDS-Rev are effective on flow-matching models, regardless of whether Reflow fine-tuning is applied (Fig. 34, Table. 1). 2), Reflow fine-tuning enhances iRFDS performance on image inversion and editing tasks (Fig. 5, Table 2). 3). When applied to score-matching-based methods like diffusion models, RFDS is equivalent to the SDS loss. RFDS-Rev offers slight improvements over the SDS baseline. While iRFDS may not produce the best editing results with diffusion models, it presents a simple alternative to other inversion methods.
Additional advantages over diffusion based priors. Besides the performance improvements, we observe that methods based on rectified flow achieve faster convergence during the optimization of 3D models. Intuitively, if we rearrange Eq. 8, $x$ can be regarded as learning toward a direct objective $v_{\phi}+\epsilon$. In contrast, diffusion models rely on the difference in predicted noises to guide the optimization process. A detailed visual comparison is included in Appendix Sec F.
@@ 6 Conclusion, Limitations and Broader Impacts
Limitations. Since the 2D model lacks training with camera pose information, our proposed methods encounter similar issues to the SDS and VSD losses in 3D generation, such as the multi-face issue. These limitations can be addressed in future studies by training pose-aware models using multi-view data, similar to current diffusion based approaches (Ye et al., 2023; Shi et al., 2023a; Li et al., 2023; Shi et al., 2023b; Long et al., 2023). Moreover, currently our proposed iRFDS method does not taken the CFG mismatch issue into considerations. A promising future direction will be combining the iRFDS method with null-inversion (Mokady et al., 2023) to address the CFG issue.
@@
Broader Impacts. Our approach significantly expands the potential applications of pretrained rectified flow models. We demonstrate that a text-to-image model can be adapted for both image editing and text-to-3D generation tasks. However, since we utilize pretrained text-to-image rectified flow models as our foundational network, our methods might also inherit the biases present in these base networks.
Appendix
A Proof: RFDS, iRFDS, RFDS-Rev with Diffusion models
RFDS is Identical to SDS When Expressed in Terms of Score Function
As proven in Stochastic Interpolants (Albergo et al., 2023), the velocity $\bm{v}(\bm{x_{t}})$ can be expressed in terms of a score function $\bm{s}$ learned with score-matching objective
$\bm{v}(\bm{x_{t}})=\frac{\sigma_{t}\bm{s}(\bm{x_{t}})(\dot{\alpha_{t}}\sigma_{t}-\alpha_{t}\dot{\sigma_{t}})+\alpha_{t}\bm{x_{t}}}{\alpha_{t}}.\;$ (11)
For the detailed proof of this relation, we refer the readers to Albergo et al. (2023).
By substituting this relation into RFDS (Eq. 8) and considering the relation $\bm{s}=\frac{-\bm{\epsilon_{\phi}}}{\sigma_{t}}$, we directly obtain the same equation as the SDS loss
$\nabla_{\bm{\theta}}\mathcal{L}_{\text{rfds}}(\bm{\phi},\bm{x},\bm{\epsilon},t)\simeq\mathbb{E}\left[w(t)\underbrace{(\bm{\epsilon_{\phi}}(\bm{x_{t}})-\bm{\epsilon})}_{\text{Score Residual}}\underbrace{\frac{\partial\bm{x}}{\partial\bm{\theta}}}_{\text{Generator Jacobian}}\right].$ (12)
iRFDS Expressed in Terms of Score Function
Similarly, we can derive the iRFDS in terms of score function
$\nabla_{\bm{\epsilon}}\mathcal{L}_{\text{irfds}}(\bm{\phi},\bm{x},\bm{\epsilon},t)\simeq\mathbb{E}\left[w^{\prime}(t)\underbrace{(\bm{\epsilon_{\phi}}(\bm{x_{t}})-\bm{\epsilon})}_{\text{Score Residual}}\right].$ (13)
After the two formulas are derived, we can naturally arrive at RFDS-Rev.
B Implementation Details
@@
RFDS and RFDS-Rev for 3D generation. We use the 3D model implicit model Instant-NGP (Müller et al., 2022) as the 3D backbone. Each 3D model is optimized for 15000 steps. We use a CFG of 50 for all 3D experiments and 2D toy experiments. The model is optimized with a resolution of 256 for the first 5000 steps and then 500 for the final 10000 steps. The experiments are carried out on NVIDIA A6000 GPUs. On InstaFlow, generating one 3D scene using RFDS-Rev takes around 30 mins wall-clock time. On SD3, generating one 3D scene using RFDS-Rev takes around 1 hour. We use $w(t)=1$ on SD3 and $w(t)=-1$ on InstaFlow. Choosing iRFDS step size is important to achieve high quality generation with RFDS-Rev. For InstaFlow, we use a stepsize 1. For SD3, we observe that a stepsize of $1-\sigma_{t}$ produce reasonable results.
@@
iRFDS for image inversion and editing. The inversion starts from a randomly sampled Gaussian noise. We optimize the noise for 1000 steps using iRFDS and CFG 1 with a learning rate of 3*10-3. To facilitate effective noise optimization, we add one additional loss to enforce the noise follows a Gaussian distribution. Specifically, we add a loss to enforce the mean and variance of the current noise to be zero and one respectively. After the noise is optimized, we change the caption to the target caption and run the forward flow for 5 steps using CFG 1.5. We use $w^{\prime}(t)=-1$ on SD3 and $w^{\prime}(t)=1$ on InstaFlow.
C Additional Related Works